# Using machine learning to measure evidence of students’ sensemaking in physics courses

### v2

@@ Abstract
In the education system, problem-solving correctness is often inappropriately conflated with student learning. Advances in both Physics Education Research (PER) and Machine Learning (ML) provide the initial tools to develop a more meaningful and efficient measurement scheme for whether physics students are engaging in sensemaking: a learning process of figuring out the “how” and “why” for a particular phenomena. In this work, we contribute such a measurement scheme, which quantifies the evidence of students’ physical sensemaking given their written explanations for their solutions to physics problems. We outline how the proposed human annotation scheme can be automated into a deployable ML model using language encoders and shared probabilistic classifiers. The procedure is scalable for a large number of problems and students. We implement three unique language encoders with logistic regression, and provide a deployability analysis on real student explanations from the 2023 Introduction to Physics course at Tufts University. Furthermore, we compute sensemaking scores for all students, and analyze these measurements alongside their corresponding problem-solving accuracies. We find no linear relationship between these two variables, supporting the hypothesis that one is not a reliable proxy for the other. We discuss how sensemaking scores can be used alongside problem-solving accuracies to provide a more nuanced snapshot of student performance in physics class.
@@ I Introduction
Teaching and instruction in undergraduate physics courses has largely relied on problem-solving as the standard method to measure student performance [1, 2, 3, 4, 5, 6]. Common practice is for “real-time” performance to be measured via multiple-choice or single-solution problems, where canonically correct answers determine the student’s knowledge of the core material. Accuracy scores across assignments and examinations, typically coupled with letter grades, act as signals of progress throughout the course as well as final verdicts of student success.
While engaging in problem-solving is a useful experience for students in a physics classroom, using the problem solution as a measure of student learning assumes a direct correlation that may not always hold. Problem-solving accuracy as a measurand assumes that students will engage in a learning process involving the core material to obtain the problem solution. Often times, there are alternative strategies for obtaining a problem solution such as rote-memorization of the rules or procedures required for solving similar problem types [7]. In this scenario, students would score very high on exams that contain these problem types; however given a previously unseen problem structure where the same core material is to be applied, the students would struggle. Here, a risk of using problem-solving accuracy as the predominant metric is an inflated sense of confidence in both the instructor and the student that the core material has been learned. It could also pose a risk for confounding variables in research studies that aim to investigate how instructional techniques influence student learning [8, 9, 10, 11, 12].
Over the last fifteen years, researchers in STEM Education (Ed.) have focused more closely on modeling a student’s learning process directly - a process typically referred to as sensemaking [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Different from memorizing facts, procedures, or problem types, sensemaking is a macro-cognitive process that allows one to construct a model of “how” and “why” with respect to the core material that is consistent with prior experiences and integrates new ones [13]. In physics courses, there is an emphasis on physical sensemaking, where the goal is for students to construct models that describe physical phenomena. This is separate from (blended) mathematical sensemaking, which focuses on how students understand equations and symbolic forms as physical concepts across various scientific disciplines [26, 27, 28, 29, 30, 31, 32, 33, 34].
Prior work has primarily focused on characterizing physical sensemaking through numerous qualitative analyses of in-moment social dynamics and individual student interviews [25, 23, 14, 15, 16, 17, 18, 19, 20]. In these scenarios, students were given a problem to solve, but the research objective was focused on modeling the sensemaking process that took place as a problem-solving strategy. These empirical investigations resulted in recognizable qualitative outcomes of the physical sensemaking process. It is thus a natural next step to consider whether these outcomes can be turned into a post-hoc measurement scheme that reliably quantifies evidence of physical sensemaking given student explanations. Crucially, we need such a scheme that scales well with student population size.
In this paper, we take this next step forward and answer three questions: (1) How can we measure evidence of students’ post-hoc physical sensemaking across multiple problems? (2) How can we use machine learning tools to automate this measurement procedure? (3) How can these measurements be used to obtain a more informed picture of student performance in a physics course?
Our main theoretical contribution is a measurement scheme, which quantifies the evidence of students’ physical sensemaking post-hoc by annotating their written explanations for their solutions to physics problems. We outline how this human annotation scheme can be transformed into a deployable machine learning (ML) model using language encoders and shared probabilistic classifiers, freeing up time for a human annotator. We implement three ML model variations with different text encoders, and provide a deployability analysis on real student explanations to four problems from the 2023 Introduction to Physics course at Tufts University. A pipeline based on the BERT language encoder [35] that was fine-tuned for our task achieves the best Area Under the Receiving Operator Characteristic (AUROC) score of on the test set, which contains explanations from students not seen in training data on the same problems as in the training data. Lastly, we compute sensemaking scores for all students, and analyze these measurements alongside their corresponding problem-solving accuracies. The computed Point-Biserial correlation coefficients between the two variables signify that we cannot claim there exists a linear relationship between problem-solving accuracy and evidence of physical sensemaking in the classroom. This demonstrates a need to measure evidence of the sensemaking process explicitly, rather than relying on problem-solving accuracy as a proxy.
Alongside this manuscript, we contribute practical materials that can be used to help other individuals measure physical sensemaking: (1) examples of implementing the measurement scheme on four physics problems with the corresponding human annotated dataset (385 total student explanations) and (2) ready-to-implement code that generates physical sensemaking measurements for new student explanations on the same set of four problems. Both of these materials can be found in the Github Repository “AuSeM” 111Github repo: https://github.com/tufts-ml/AuSeM.git.
II How to Measure Post-Hoc Physical Sensemaking
II.1 Defining Physical Sensemaking
Given a phenomena of interest, physical sensemaking is the macro-cognitive process that constructs the mental model of “how” and “why” for a given physical phenomena by integrating prior and new experiences until consistency is achieved [13]. Inspired by prior works across research fields that have attempted to theoretically model and empirically characterize student sensemaking [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 37] we summarize the key features of this process: disassembling the phenomena into components, identifying characteristics (i.e. feature observations), and considering both descriptive (i.e. associations) and mechanistic (i.e. cause and effect) relationships. During the process, a student may test various hypotheses to “figure things out.” When someone has sensemade, they have formed an organized mental model of these components, characteristics, and relationships that is self-consistent. The phenomena can then be explained with natural and mathematical language. These prior works, which we discuss in greater detail in Sec. II.4, provide the foundational starting point for a quantitative post-hoc measurement scheme. Let us consider a problem which contains a physical phenomena to be sensemade as a part of the solution strategy. Provided a student’s text explanation of how they solved the problem, a human annotator can ask themselves the following questions:
@@ II.1 Defining Physical Sensemaking
Does the student identify the components that are relevant to the phenomena?
@@
Does the student identify the characteristics of the components that are relevant to the phenomena?
@@
Does the student identify the descriptive and mechanistic relationships that are relevant the phenomena?
These initial questions are general and can be further divided into criteria questions that capture more specific outcomes of the sensemaking process. Thus, we propose 8 domains that represent what we expect to observe in the student’s explanation as a result of physical sensemaking.
1. Objects. The central entities that interact to form the physical phenomena in a single hierarchical level. Each object object interaction is relevant to the phenomena. Examples: “car”, “box”, “ball”, “people”, “particles”, etc.
2. Influences. The non-central environmental entities that interact with objects. Only the influence object interaction is relevant to the phenomena. Examples: “Earth’s gravitational field”, “human interference”, “electromagnetic field”.
@@
3. Properties. The relevant qualities that are confined to the physical boundary of an object. Examples: mass, temperature, volume.
@@
4. Positioning. The spacial states or orientations of an object. Examples: “downwards”, “upwards”, “left”, “right”, “x-direction”, “y-direction”.
@@
5. Movements. The motion states of an object. Examples: “initial velocity”, “slowing down”, “speeding up”.
@@
6. Interactions. The exchanges between all object object and influence object pairs. Examples: “gravitational force”, “human pushing”, “electromagnetic force”.
7. Descriptive Relationships. The associations between properties, positioning, and movements of an object. Examples: “a change in position is a velocity of an object”, “a change in temperature is associated with a change in velocity of a particle”.
@@
8. Mechanistic Relationships. The changes that result from object object interactions and influence object interactions. Examples: “constant acceleration of a ball caused by Earth’s gravitational field exerting a downward net force”; “acceleration of a book initially at rest caused by a human push”.
Provided a student’s text explanation as to how they arrived at a particular solution for a problem, the annotator can ask themselves whether the student identifies criteria in each of the 8 domains. For example, see Figure 1. This is a student explanation in response to Problem 4 [38] below:
Example: Problem 4 1. The net force on the box is in the positive x direction. Which of the following statements best describes the motion of the box? (a) Its acceleration is parallel to the x axis. (b) Its velocity is parallel to the x axis. (c) Both its velocity and acceleration are parallel to the x axis. (d) Neither its velocity not its acceleration need to be parallel to the x axis. 2. Briefly explain your answer to the previous question.
The physical phenomena for a student to sensemake in this problem is an object’s motion under a net external influence. Note that our measurement scheme is only made possible by the student’s response to part 2, while part 1 only captures problem-solving correctness. We interpret the response as a student’s expression of their problem-solving strategy, and we use it to measure evidence of their physical sensemaking (as shown in Fig 1).
II.2 Post-Hoc Measurement as Multi-Label Binary Annotation
Figure 1 demonstrates how an individual student explanation can be annotated with respect to 8 domains; however note that not all 8 domains are relevant (in Fig 1 the descriptive relationships domain is not relevant to the particular problem) and some domains occur multiple times (e.g. the object has more than one positioning in the context of the overall problem). As such, the full set of yes/no criteria questions across the 8 domains is problem-specific. Thus, for a set of problems, an annotator can follow the proposed steps:

### v3

# Combining physics education and machine learning research to measure evidence of students’ mechanistic sensemaking

@@ Abstract
Advances in machine learning (ML) offer new possibilities for science education research. We report on early progress in the design of an ML-based tool to analyze students’ mechanistic sensemaking, working from a coding scheme that is aligned with previous work in physics education research (PER) and amenable to recently developed ML classification strategies using language encoders. We describe pilot tests of the tool, in three versions with different language encoders, to analyze sensemaking evident in college students’ written responses to brief conceptual questions. The results show, first, that the tool’s measurements of sensemaking can achieve useful agreement with a human coder, and, second, that encoder design choices entail a tradeoff between accuracy and computational expense. We discuss the promise and limitations of this approach, providing examples as to how this measurement scheme may serve PER in the future. We conclude with reflections on the use of ML to support PER research, with cautious optimism for strategies of co-design between PER and ML.
@@ I Introduction
Advances in machine learning (ML) and natural language processing [1, 2, 3, 4, 5] have generated a rush to develop applications in science education, often focused on supporting the analysis and assessment of student thinking [6]. Much of the promise is to make coding or grading of student work more efficient, as manual assessment of student thinking requires time-intensive, subjective human judgment. A key challenge is to provide successful ML support without burdensome requirements for expensive computational or data resources.
At the same time, scholars argue for patience and care: AI does not provide a ”magic bullet” [7] for research or instruction. The benefits of these new technologies will depend on careful investigation, including collaboration and co-design between ML and physics education researchers. Examples include recent work by Fussell and her colleagues to propose methodology for assessing the trustworthiness of machine coding, illustrated with analyses of students’ open responses to questions about empirical science, [8]. In subsequent work, they compare results across different large language models (LLMs) in analyzing skills evident in student lab notes [9].
We are focusing on the challenge of measuring the students’ sensemaking, which researchers have identified as an essential target for research and instruction, but have been constrained to study in specific episodes or small sets of data [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. These studies took place in a variety of situations–different subject matter, levels of education, forms of data–and there is substantial variation across them in meanings of and nature of evidence for sensemaking. Odden and Russ [10] described the literature as “theoretically fragmented.” To resolve this issue, they proposed a definition of sensemaking as “a dynamic process of building or revising an explanation in order to ‘figure something out’ – to ascertain the mechanism underlying a phenomenon in order to resolve a gap or inconsistency in one’s understanding.” We take this as a starting point for our work, recognizing several features we intend ultimately to study: (1) the perception of some gap or inconsistency in one’s understanding; (2) reasoning with the purpose of addressing that gap; and, for science broadly and physics in particular, (3) an expectation of a mechanism underlying the phenomenon [22].
For now, we focus on developing ML to analyze student reasoning with respect to mechanism, drawing on an earlier analytic framework by Russ and her colleagues for human analysis of mechanistic reasoning [22]. We review that framework below and discuss the need for its adaptation. Designing an approach to ML involves a myriad of contingencies and decisions that depend on the particularities of the situation. Here, the particular data we study consist of college students’ brief 1-5 sentence explanations for answers to multiple choice conceptual questions in introductory mechanics. The data comes from from a 100-student calculus-based course at Tufts University, which used FlipIt Physics [23] and emphasized students’ learning to engage productively with confusion [24]. Given this data, two specific questions motivate our current efforts. RQ1: How can we refine an analytic framework from physics education research (PER) into a measurement scheme amenable to ML with this data set? RQ2: For ML implementations of this scheme, how do design decisions that limit the computational resources needed impact how well the ML-provided labels agree with human coding?
The remainder of this article is organized as follows. Sec. II provides background with respect to the data we hope to analyze, prior work on mechanistic reasoning, and prior work to incorporate ML in science education research. Sec. III presents an account of adapting the framework from Russ et al. [22] into a measurement scheme for our particular data set, addressing RQ1. Note that we will use the term scheme and the phrase measurement scheme throughout the paper to refer to this adaptation, which is amenable to ML. We present the scheme with examples of its application, and we describe the ML in technical detail. Part of our purpose in this account is to provide an example of convergence between PER and ML in the co-design of analytic tools.
Sec. IV delves into a specific aspect of the design, RQ2, to compare ML performance across three possible language encoders. In this section, we consider both the computational resources required for the model and the level of agreement with a human coder. We describe how this approach, which currently depends on problem-specific training, may eventually generalize to support analysis of student responses to new problems without that step of training. In Sec. V, we discuss the promises and limitations of this use of ML, specifically touching on utility, efficiency, generalizability, reliability, sustainability, and privacy.
Finally, in Sec. VI, we reflect on our efforts to date, considering both our specific aims in developing this particular tool as well as our and the broader community’s more general aims of developing artificial intelligence for applications in education. As will be clear, we are not presenting settled science. Rather, we are giving examples to illustrate challenges and possibilities of this work.
To enable others’ explorations, we provide open-access data and code resources in a public Github repository linked at the bottom of page one, under the project name “AuSeM”. There, we release our dataset of 385 total student responses across 4 physics problems, each manually-annotated using our approach. We also provide a codebook of instructions for human annotators, which reflects updated clarifications that emerged from our inter-rater agreement process. These resources could be used to train other models or as a guide for further manual annotation efforts of other data. Our provided Python code can be used to reproduce our experiments (re-train each architecture on our data) or generate mechanistic sensemaking scores for new student responses on the same set of four problems.
II Background
II.1 The challenge to analyze sensemaking
The context for this effort is a calculus-based introductory course in Newtonian Mechanics at Tufts University. The course places unusual emphasis on students’ sensemaking, with explicit goals of their learning to recognize and to engage proactively with confusion and uncertainty [25]. The data we examine come from spring 2022, when the course was using FlipIt Physics [23]. Before each class all semester long, students watched a 20-minute “prelecture” video introducing new concepts, in preparation for the in-person lecture class. After the video, students were required to answer a multiple-choice conceptual question to check their understanding. For each problem, they submitted an answer as well as a written text explanation of their choice.
As a motivating example, consider the following problem from the fourth unit of the curriculum:
Example: Problem 4 1. The net force on the box is in the positive x direction. Which of the following statements best describes the motion of the box? (a) Its velocity is parallel to the x axis. (b) Its acceleration is parallel to the x axis. (c) Both its velocity and acceleration are parallel to the x axis. (d) Neither its velocity not its acceleration need to be parallel to the x axis. 2. Briefly explain your answer to the previous question.
Our challenge is to analyze written text data from student explanations as evidence of their engagement in mechanistic sensemaking. Here are some examples of students’ explanations:
E1: “According to Newton’s second Law, F(net) = ma, we could know that the direction of a is the direction of F(net)” (Answer b)
E2: “If the force is acting on it in the positive x direction, then the box must be moving in the positive x direction and will therefore have a parallel acceleration to the x axis since it is not speeding up. The velocity would also be parallel because its direction is positive. I disagree with myself as I type this but I’m not sure what my opposing argument is.” (Answer c)
E3: “If a 3 people were pushing the box but one person was pushing harder than the rest, the box would accelerate in the same direction as the greater force. However, with velocity, the final velocity has to take into account the initial velocity. Though the acceleration might go to the right, if the velocity was going upwards, the box might move at a slant.” (Answer b)
It is a challenge to decide what to take as evidence of sensemaking, or of not-sensemaking. The first makes sense to a physicist, but is it evidence of sensemaking for the student? The second does not make sense to a physicist (the box accelerates “since it is not speeding up”?), but the student noted experiencing uncertainty; perhaps it is evidence of their trying to identify that uncertainty. Example E3 makes sense to a physicist, and it also shows the student constructing a tangible physical situation.
Ultimately, we hope to make progress in tapping these data for analysis of sensemaking, to study whether and how students in the course make progress in learning how to learn. In this work, we start in that direction by analyzing for evidence of mechanistic reasoning. Mechanism is consistently described in accounts of sensemaking in physics [19, 21], and there is a well-established framework for analysis [22].
II.2 Mechanistic reasoning in science
Russ and her colleagues [22] adapted work by philosophers of science Machamer, Darden, and Craver [26], who had analyzed the concept of mechanism within areas of biochemistry. This led to an analytic framework made up primarily of seven levels organized in a hierarchy:
#1. Describing the target phenomenon that is observable and reproducible.
#2. Identifying setup conditions in the spatial and temporal organization of the entities that lead to the phenomenon.
#3. Identifying entities involved in producing the phenomenon.
#4. Identifying activities, the “actions and interactions that occur among the entities.”
#5. Identifying properties of entities that are relevant for producing the phenomenon.
#6. Identifying the spatial organization of entities that are specifically relevant for the actions and interactions among the entities.
#7. Chaining backward or forward, in inferences about the entities and activities that could have led to (backward) or could follow from (forward) a given situation.
These seven categories are logically hierarchical in that higher levels generally require lower levels. For example, chaining forward or backward would generally require identifying entities, properties, and activities. Russ et al. [22] added two categories they considered non-hierarchical:
Analogies as explicit comparisons to other phenomena. Animated models in students’ use of physical actions, whether manipulating objects available or gestures, to simulate the entities and activities they are describing.
In the 20 years since its introduction, this framework has been applied broadly in PER and other areas of science and engineering education research [27, 28, 21].
The first example explanation from our data, labeled E1 above, does not include any entities or activities; there is no evidence of mechanistic reasoning. That does not mean there was no mechanistic reasoning; only that none is evident in the provided text. Nor does it mean the reasoning is bad; it is correct.
The second explanation E2 gives clear evidence of entities (the box) and activities (its movement). Perhaps the student is also thinking of the force as an entity? Perhaps, too, there is a chaining, a force causing motion, but the logic is difficult to follow–including evidently for the student. It is “moderate evidence of mechanistic reasoning,” by Russ et al’s framework.
Example E3 has evidence of entities (people, the box), activities (pushing, motion), and of chaining (the box would accelerate in the direction of the greater force/push). There is also evidence, although it is not quite as clear, of spatial organization (if the force is sideways, by the people pushing, but the box is moving upward, the box might move at a slant). Overall, this example is “strong evidence of mechanistic reasoning.”
Note that applying Russ et al.’s 7-level general framework to this problem involves both specification and simplification. In our data, student responses do not describe a target phenomenon, as the problem prompt does that already. Within the identification steps, there are only a small set of possibilities for what could arise as entities and activities. As we describe below, in our work to date we take this step, seeking a measurement scheme that can be applied both manually and via ML.
II.3 Incorporating ML into science education research
ML tools are now rapidly integrating into PER, predominantly via text classification methods - algorithms that learn patterns in student text responses and sort each response into distinct categories or labels [29, 30, 31, 32, 33, 34]. This rapid integration has been largely driven by recent advances in large language models (LLMs) [1, 2, 35] trained on vast amounts of language corpora (books, Wikipedia, websites, etc). The availability of chatbots like ChatGPT powered by LLMs makes it possible for PER researchers to directly feed text into a chatbot window to be classified [36]. Below, we contrast different strategies for using LLMs in PER.
One paradigm for LLM-enabled PER is in-context learning (ICL), where an existing chatbot is provided a student text response and prompted via natural language instructions to classify it into a desired labeling system. To be clear, the “learning” here is euphemistic; ICL does not entail retraining the LLM itself at all. Instead, it relies on the pretrained ability of the LLM to transform an instructional prompt and provided input text into useful output text. ICL is data-efficient, with proponents often providing only a few example demonstrations in the instructions to clarify the desired labeling transformation. However, ICL has a number of important limitations. First, ICL’s utility relies on the assumption that the LLM is useful for a specific downstream language task in education (correctness, reasoning, sensemaking, etc.). The underlying LLM was not trained to do well on these tasks, and it is important to verify at scale whether an LLM is reliable at certain tasks. ML researchers have sometimes questioned the utility of ICL, showing that “demonstrations have a marginal impact” [37] and “randomly replacing labels in the demonstrations barely hurts performance” [38]. Second, ICL is sensitive and variable. Outputs can vary depending on the exact prompt [39] and the selection and order of examples [40]. Even the same exact input may lead to different outputs due to the stochastic nature of chatbots. Third, LLMs reflect biases inherent in the source data [41]. Finally, there is a lack of transparency: closed-source versions such as ChatGPT do not allow users to inspect, understand, or modify how outputs are obtained. For these reason, in-context learning methods deserve and require further investigation for their reliability in science education research.
Other works focus on training supervised ML predictors to accomplish their specific task, integrating accepted human annotation frameworks in science education with machine learning practices [31, 32, 28]. Watts et al. [28] is a prime example of this type of integration between science education and machine learning research. Watts et al. [27] had previously adapted Russ et al. [22] for use in organic chemistry; follow-up work [28] adapts further to apply ML binary classifiers. This entailed separating Russ’s original categories into stand-alone binary (present or not present) codes specific to organic chemistry problems. For example, they defined the specific code electron movement within Russ’s identifying activities. The authors treat each code as independent evidence of mechanistic reasoning, without the hierarchal organization feature of Russ’s framework, which simplifies the ML methodology, a useful first step towards integrating ML and science education research. Student responses are used as training data for separate binary classifiers for each code, where various classifier types are attempted, and only the final results of the highest performing classifier are reported.
Similar to Watts et al, we take the step of adapting the framework to specific features of the tasks in our data set, and we simplify to binary codes, as we describe in the following section. We will discuss similarities and differences in technical detail in Sec. III.3.
Fussell et al. [8] provide an evaluation of different large-language models for measuring specific skills in students’ typed lab notes. The authors compare off-the-shelf pre-trained open-source LLMs with ones that are fine-tuned (further trained) on training data specific to their task. The results showed that higher-resource models (e.g. more training data, computational power) often outperform lower-resource models for their particular task, although depending on the evaluation metric there are exceptions. More specifically, the authors found that all fine-tuned models outperformed the off-the-shelf pre-trained models, and they suggest that for theory-driven classification tasks (i.e. where there is a specific task in science education), fine-tuning might be more useful.
There are a few key distinctions between our current work and Fussell et al. Similarly to Fussell et al., we compare the performances of ML model variants in the context of our specific task; however, we do not attempt to provide more general insights about the models on other tasks or datasets. We will discuss the similarities and differences between Fussell et al. and our work in more technical detail in in Sec. III.3.
III RQ1: PER-ML Measurement Scheme
III.1 Refining the framework for a particular data set
Russ et al. [22] designed their framework to be applicable across areas and levels of scientific reasoning. That breadth of applicability has been helpful for the research community, but its generality presents challenges for the development of an ML tool with context-specific datasets. Given a particular data set, such as the conceptual question above about a net force acting on a box, it is possible and advantageous to tune it to the specifics of the task. The strengths and limitations of machine learning motivate other simplifications as well.
Task-specific considerations. We are interested in measuring student responses to how they solved a particular problem within a set of closed-ended introductory physics problems. Each problem contains its own specific entities, properties, activities, etc. that are relevant. The box question, for example, mentions only one entity, and there is a limited set of entities (such as people to push) that students might add to the situation. As such, the mechanistic evidence that we are looking for is problem-dependent, which allows us to simplify the scheme in several ways.
(1) No need to re-describe the target phenomenon: Each problem’s instruction text makes the target phenomenon clear; students focused on brief explanations and thus did not restate that phenomenon in their own words. For these reasons, we do not include this first code from Russ’s framework in our scheme.
(2) Restricted set of entities, properties, and activities: The “net force on the box” example problem names only the box as an entity. There are in principle other entities one could mention as involved, such as the Earth or the surface under the box, but no students in this data set mention them. For each problem in our data, we confine our analysis to a limited, analyst-defined set of entities, properties, and activities that could reasonably be relevant or that arise in the data.
ML design considerations. While we want custom annotations for each problem, developing a separate classifier for each problem would be prohibitively expensive and unlikely to work well without many training examples. We are thus interested in combining powerful language encoders [1, 3, 35, 2] with shared classification heads that take advantage of similar word occurrence and grammatical structure across problems to efficiently annotate new examples. As such, it is important to construct multi-label classification categories that share similar linguistic features across student responses and problems. This influences three key design decisions.
(1) Share initial and subsequent spatial organizations: We assume a description of the spatial organization of entities would contain similar semantic and structural features independent of whether it refers to the initial, intermediate, or final conditions. These temporal states can be represented by a separate context embedding (further discussed in Sec. III.3), such that evidence for each temporal state can still be individually annotated, but spatial linguistic features can be shared.
(2) Separate central from non-central entities: By definition, central entities are connected to the other codes (properties, activities, etc.). As such, central and non-central entities will contain dissimilar relationships among words in a student response. Keeping these categories separate is especially useful for future work that incorporates dependencies between domain labels.
(3) Separate movement activities from interactions: Descriptions of individual entity movement activities (e.g. the object is moving with an initial velocity) are different from descriptions of the interactions between entities (e.g. the hand pushes the book) both semantically and structurally within a sentence.
Like Watts et al. [28], we relax the hierarchy of Russ’s framework, coding only for the presence or absence of evidence in each category. This simplifies our annotation scheme and ML method for now; later in Sec. VI we discuss plans going forward to reintroduce some form of that hierarchy. This would be an improvement on our current modeling assumptions: for example, evidence of properties is likely to correlate with evidence of entities.
Proposed domains. Given the above considerations, we propose 7 domains, broad categories of yes/no questions that measure evidence of mechanistic sensemaking in student thinking via the framework:
@@ III.1 Refining the framework for a particular data set
Objects. Identifying the central entities that interact to comprise the physical phenomena in a single hierarchical level. These objects have properties, spatial layout, and movement relevant to the solution process. Examples: “car”, “box”, “ball”, “particles”, etc.
@@
Influences. Identifying the non-central environmental entities that affect the objects and their activities. Only the influence object interaction is relevant to the phenomena. Examples: “Earth’s gravitational field”, “human hand”, “electromagnetic field”.
@@
Properties. Identifying the relevant qualities that are confined to the physical boundary of an object. Examples: mass, temperature, volume.
4.
@@
Spatial organization. Identifying the spatial states or orientations of an object. Examples: “downwards”, “upwards”, “left”, “right”, “x-direction”, “y-direction”.
5.
@@
Movement activities. Identifying the motion states of an object. Examples: “initial velocity”, “slowing”, “speeding”, “accelerating”.
6.
@@
Interactions. Identifying the physical exchanges between all object object and influence object pairs. Examples: “gravitational force”, “human pushing”, “electromagnetic force”.
7.
@@
Chaining. Identifying the characteristic (a property, spatial organization, or movement activity) changes that result from specific object object interactions and influence object interactions. Examples: “constant acceleration of a ball caused by Earth’s gravitational field exerting a downward net force”; “acceleration of a book initially at rest caused by a human push”.
Our scheme corresponds with Russ’s framework as follows: the properties domain directly maps to Russ’s category #5; the objects domain and influences domain combine to obtain #3; the spatial organization domain encompasses #2 and #6; the movement activities and interactions domains combine to form #4, and the chaining domain maps to #7. We do not consider stating the target phenomena, analogies, or mental model animation in this work - although these are interesting additions for future research.
Example problem and associated coding. Consider again the “force on a box” problem from Sec. II.1. An example student explanation is shown in Fig. 1, with specific parts color-coded by our 7 domains. This coding demonstrates how parts of an individual student response can be associated to each of the 7 domains. These specific parts could be later used to answer yes/no questions to quantify evidence of sensemaking, as described in Sec. III.2. Some domains correspond to multiple parts of the response (e.g. the object has more than one spatial organization in the context of the overall problem). Note that not all 7 domains need to be coded for if some are irrelevant to a problem, though all domains are relevant to the example problem here.
III.2 Measurement as multi-label binary annotation
Given a specific physics problem, our measurement scheme requires the annotator to develop a list of yes/no criteria questions for each of the 7 domains that is problem-specific. Thus, given a dataset of potential problems and student responses to be measured, an annotator can follow the proposed steps: